How do you simplify the following:
1. 3rd order radical of 432x^8 (instead of square root it is the cube root)
2. Solve: Square Root (2x+6)= 2 Square Root x
Thank you

Answer by stanbon(57984)

#1 (432*8)^(1/3)= 8^(1/3)[8*27]^(1/3)=8^(1/3)*8^(1/3)*7^(1/3)=2*2*3=12
#2 sqrt(2x+6)=2sqrt(x)
Square both sides to get:
2x+6=4(x)
Subtract 2x from both sides to get:
2x=6
Divide by "2" to get:
x=3
Cheers,
stan H.